Find T1 and T2. If the tension each rope can withstand is 800N, will they survive?

Mi = 175 kg M2 = 154 kg Ti I wish I paid attention in physics T2 + M board = 15.0 kg 2.00 m 3.00 m 8.00 m
Answer #1

FBD of plank om M28 Mg Mboardy im bom Balancing torque about o Tx8= (1.7x6) 4(M...y x4) ( 3x3) T. X8(17586)+(15x4) + (154x3)J1925.7 + T2 = (175 +15+154)x9.8 = 33712 T2 = (3371.2 - 1925.7) N = 1445.5 NT1 and T2 both are greater than 800 N .so they will not survive.